# An Economical Approach to Design with Precision Criteria

### v3

@@ Abstract
Estimation frameworks for statistical inference are preferred to hypothesis testing when quantifying uncertainty and precisely estimating effect sizes are more valuable than binary decisions about statistical significance. Study design for estimation-based investigations often uses precision criteria to select sample sizes that control the length of interval estimates with respect to a sampling distribution. In this paper, we formally define a distribution that characterizes the probability of obtaining a sufficiently narrow interval estimate as a function of the sample size. This distribution can be used to determine the smallest sample size needed to ensure an interval estimate is sufficiently narrow. We prove that this distribution is approximately normal in large-sample settings for many data generation processes. However, this approximate normality may not hold for studies with moderate sample sizes, particularly when incorporating prior information or obtaining asymmetric interval estimates. Thus, we also propose an efficient simulation-based approach to approximate the distribution for the sample size by estimating the sampling distribution of interval estimate lengths at only two sample sizes. Our methodology provides a unified framework for design with precision criteria in Bayesian and frequentist settings. We illustrate the broad applicability of this framework with several examples.
@@ 1 Introduction
Practitioners across disciplines design studies to collect data that meaningfully inform decision making in the presence of uncertainty. Different study objectives – such as assessing the plausibility of hypotheses or precisely estimating quantities of interest – motivate distinct approaches to study design. While many studies are designed with the goal of rejecting a null hypothesis, the null hypothesis significance testing (NHST) framework has been scrutinized for its association with the scientific replicability crisis (Open Science Collaboration, 2015). The American Statistical Association replied to these concerns by releasing an official statement on $p$-values (Wasserstein and Lazar, 2016) that challenged the statistical community to develop alternatives and extensions to the traditional hypothesis testing framework. This statement prompted various responses including a special issue of The American Statistician that discussed “moving to a world beyond p<0.05p<0.05(Wasserstein et al., 2019), a call to action from the President of the Institute of Mathematical Statistics (Meng, 2019), and a statement published in the Annals of Applied Statistics on statistical significance (Benjamini et al., 2021).
@@
To accommodate Bayesian and frequentist approaches, we broadly introduce precision criteria for interval estimates. We henceforth use the term CI for statements that apply to both confidence intervals and credible intervals. Our methods accommodate various types of CIs – including frequentist confidence intervals as well as equal-tailed credible intervals or those defined using regions of highest posterior density. For now, we denote the length of the $100\times(1-\alpha)\%$ CI for the target of inference $\theta$ as $L_{1-\alpha}$. We later emphasize that this CI is a function of the data. Precision criteria generally select a sample size $n$ to ensure a large probability that $L_{1-\alpha}$ is sufficiently narrow with respect to the sampling distribution for the future data. The length probability criterion (LPC) was defined to formalize this objective in Bayesian settings (De Santis and Pacifico, 2004; Brutti et al., 2014). In a Bayesian or frequentist setting, the LPC could be applied to select the smallest sample size $n$ such that $\Pr(L_{1-\alpha}\leq l)\geq q$ for some target length $l$ and probability $q\in(0,1)$. This focus on controlling the length of CIs is not to the exclusion of the frequentist notion of coverage, which should approach the nominal level of $1-\alpha$ in frequentist and Bayesian contexts as $n\rightarrow\infty$ under standard regularity conditions (van der Vaart, 1998).
For Bayesian studies, Stevens and Hagar (2022) recently assigned a distribution to the LPC with respect to the sample size $n$ under the conditional approach. This distribution was called the sufficient sample size distribution (SSSD), and it was formally defined so that its $q$-quantile is the smallest sample size $n$ such that $L_{1-\alpha}$ will have length at most $l$ with probability at least $q$ (i.e., such that $\Pr(L_{1-\alpha}\leq l)\geq q$). The SSSD was observed to be approximately normal in a variety of situations via simulation (Stevens and Hagar, 2022), but theoretical justification for this approximate normality was not explored. Furthermore, conditions under which the SSSD exists and this approximate normality holds were not clearly defined. In this work, we define the length probability distribution (abbreviated as the LP distribution) that extends the SSSD to accommodate Bayesian and frequentist CIs under both the conditional and predictive approaches to specify sampling distributions.
@@
In this paper, we propose two methods to estimate the LP distribution across statistical paradigms. The first method analytically approximates the LP distribution using asymptotic theory. The second method leverages this asymptotic theory to inform an efficient method to estimate the LP distribution based on simulation results conducted at only two sample sizes; this method is generally applicable and flexibly accommodates moderate sample size settings where asymptotic approximations may be unsuitable. Similar approaches have recently been proposed to efficiently select sample sizes within the hypothesis testing framework (Hagar and Stevens, 2024, 2025a, 2025c, 2025b). While our design framework proposed here is theoretically intricate, its implementation is straightforward, promoting an economical and broadly useful approach to simulation-based sample size determination with precision criteria.
@@ 2 Preliminaries
Our design framework denotes data from a random, future sample as $\boldsymbol{W}^{{}_{(n)}}$, where the sample size $n$ is the number of independent sampling units (ISUs). The observed data are denoted by $\boldsymbol{w}^{{}_{(n)}}$. For non-clustered data, each observation is an ISU and in settings with clustered data, multiple dependent observations may comprise a single ISU. For each ISU, the sample $\boldsymbol{W}^{{}_{(n)}}$ may consist of response(s) $\boldsymbol{Y}$ along with a vector or matrix of additional covariates $\boldsymbol{X}$ that may encode information about treatment group assignment and/or assignment to ISUs. We assume that each ISU in $\boldsymbol{W}^{{}_{(n)}}$ is generated independently according to the model $f(\boldsymbol{w};\boldsymbol{\eta})$, where $\boldsymbol{\eta}$ denotes a vector of parameters. The target of inference $\theta$ is specified as a function $g(\cdot)$ of these parameters: $\theta=g(\boldsymbol{\eta})$. In an estimation framework, we aim to draw inference using a two-sided $100\times(1-\alpha)\%$ CI for $\theta$. We henceforth denote the length of the $100\times(1-\alpha)\%$ CI as $L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})$ for a future sample of $n$ ISUs; the length of the corresponding CI for an observed sample is $L_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}})$. This notation now emphasizes that the CI length and its sampling distribution are functions of the data.
@@
The collection of estimates $\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})\}_{r=1}^{m}$ in Line 5 of Algorithm 1 is used to estimate the sampling distribution of CI lengths under $\Psi$. The relevant $100\times(1-\alpha)\%$ CIs for $\theta$ can be estimated using any Bayesian or frequentist method that is based on likelihood functions. This flexibility accommodates the use of simple or complex interval estimates – such as approximate frequentist confidence intervals obtained using normal theory and sample summary statistics or Bayesian highest posterior density intervals based on computationally approximating the posterior distribution of $\theta$. The probability of satisfying the length criterion can be estimated as
@@ 3 Approximate Normality of the Length Probability Distribution
We prove Theorem 1 in Appendix B.2 of the supplement, but we consider its practical implications here. Theorem 1 imposes constraints on $\Lambda(\boldsymbol{\eta}_{r})$ and $\tau^{2}$ to ensure the normal distribution that approximates the LP distribution for a given length $l$ has finite mean $\mu_{l}>0$ and standard deviation $\sigma_{l}>0$. For large sample sizes $n$ corresponding to small target lengths $l$, the values for $\mu_{l}$ and $\sigma_{l}$ from Theorem 1 can be used to determine quantiles of the LP distribution under the conditional approach as stated in Corollary 1.
@@ 4 A Proxy to the Sampling Distribution of CI Lengths
Flexible, simulation-based design methods that control the length of CIs require that we estimate the sampling distribution of CI lengths for various sample sizes $n$. We approximate this sampling distribution by generating data $\boldsymbol{w}^{{}_{(n)}}$ using the straightforward process in Algorithm 1. For theoretical development, we create a proxy for the sampling distribution of the logarithm of the CI lengths. These proxies substantiate the theory that underpins our proposed methodology, but they are not used in our design methods proposed in Section 5.
@@ 6.2 Estimation of Illness Duration
We next assess the performance of our proposed methods in the context of estimating the duration of Mpox illness. The Mpox virus is predominantly spread via direct contact with the active lesions of an already infected individual (Kim et al., 2023). In previous clinical trials (Klein, 2024), Mpox illness duration was defined as the time in days between starting tecovirimat (a safe drug approved for use with smallpox) and experiencing the resolution of all Mpox lesions. For this example, we want to estimate the 0.9-quantile of this distribution for Mpox illness duration. Our subsequent numerical results serve purely illustrative purposes, but we emphasize the value in precisely estimating quantities associated with illness durations to inform public health recommendations. This example also allows us to consider settings with independent observations, account for additional covariates, and accommodate censoring.
@@
For illustration, we suppose that the resolution time $T_{i}$ and censoring time $C_{i}$ are exponentially distributed such that the exponential rates are log-linear functions of $x_{i}$:
$T_{i}\sim\text{EXP}(\gamma+\beta x_{i})~~~\text{and}~~~C_{i}\sim\text{EXP}(\xi+\rho x_{i}),$ (13)
@@
where $\gamma$, $\beta$, $\xi$, and $\rho$ are regression parameters. The to-be-observed time $Y_{i}$ for participant $i$ is the minimum of $T_{i}$ and $C_{i}$. Because the censoring process in (13) depends only on the observed covariate $x_{i}$, these censored data are considered to be missing at random. However, the Bayesian methods we use in this subsection could in theory accommodate more complex missingness mechanisms. At the design stage, we must also define a distribution for $\{X_{i}\}_{i=1}^{n}$. We suppose that $X\sim\mathcal{N}(0,\sigma_{x}^{2})$ for this example, and that the data from participants $i$ and $i^{\prime}$ are independent for $i\neq i^{\prime}$. To summarize, the parameters for our analysis model are $\boldsymbol{\eta}=(\gamma,\beta)$, but the parameters $\boldsymbol{\eta}^{+}=(\gamma,\beta,\xi,\rho,\sigma_{x})$ are used for data generation.
@@
Our target of inference used for study design is the 0.9-quantile of the distribution for the resolution time $T$ when $x=0$: $\theta=g(\boldsymbol{\eta})=\log(10)\exp(-\mu)$. We aim to estimate a $100\times(1-\alpha)\%=99\%$ posterior highest density interval (HDI) for $\theta$ having length at most $l=5$ days with probability $q=0.8$. To demonstrate the broad applicability of our methods, we design this study under the conditional and predictive approaches. For the conditional approach, we use the parameter values $\boldsymbol{\eta}^{+}_{r}=(\log(\log(10)/18),0.5,-\log(32),1,0.25)$ informed by findings from Kim et al. (2023) for all simulation repetitions $r$. These parameter values are such that the true value of $\theta_{r}$ is 18 days and lesion resolution times are censored for roughly 20% of participants. Under the predictive approach, we assume that $\gamma_{r}\sim\mathcal{N}(\log(\log(10)/18),0.07^{2})$, $\beta_{r}\sim\mathcal{N}(0.5,0.07^{2})$, $\xi_{r}\sim\mathcal{N}(-\log(32),0.07^{2})$, and $\rho_{r}\sim\mathcal{N}(1,0.07^{2})$; we note that small changes to these parameter values give rise to substantial changes in $\theta_{r}$ and the censoring proportions. We again assume that $\sigma_{x,r}=0.25$ for all repetitions under the predictive approach.
@@
To facilitate Bayesian analysis, we consider a diffuse and informative prior for $\gamma$ to explore how our methods perform. The diffuse prior for $\gamma$ is $\mathcal{N}(0,10^{2})$, and the informative prior is $\mathcal{N}(-2,0.3^{2})$. This informative prior is relatively well specified for illustration. For all simulations in this subsection, we assign a diffuse $\mathcal{N}(0,10^{2})$ prior to $\beta$. These choices result in four approach and prior combinations. For each combination, we analytically approximated the LP curve using the results from Theorem 1. For the predictive approach, this analytical approximation arises from a mixture of normal distributions that we obtained by applying Theorem 1 to 1000 simulated $\boldsymbol{\eta}^{+}_{r}$ values. We also approximated the LP curve using Algorithm 2 with $m=10^{4}$ simulation repetitions; to accommodate the predictive approach, the CI lengths were split into 10 bins based on the true $\Lambda(\boldsymbol{\eta}_{r})$ value at each sample size before constructing the linear approximations. We lastly approximated the LP curve by simulating the sampling distribution of CI lengths in increments of $n=10$. The resulting LP curve approximations are visualized in Figure 2.
@@
Figure 2: LP curves obtained using various estimation methods when the 0.9-quantile of Mpox illness duration is the target of inference. The horizontal dotted line represents $q=0.8$.
@@
Algorithm 2 again yields more accurate LP curve approximation than the analytical approximation. The performance of the analytical approximation resulting from Theorem 1 is reasonable for the settings with the diffuse priors for $\gamma$. Because the analytical approximation to the LP curve does not account for the prior distribution, it overestimates the required sample size when the relatively well-specified informative prior for $\gamma$ is used ($n=478$ vs. 456 for the conditional approach and $n=506$ vs. 484 for the predictive approach). Overall, the LP distributions under the predictive approach exhibit greater dispersion than those under the conditional approach. That is, we require less (more) data to attain the length criterion with smaller (larger) probabilities $q$. This greater dispersion in the LP distribution is a direct result of there being greater dispersion in the sampling distribution of CI lengths under the predictive approach.
@@
It took less than one second to approximate the LP distribution under the conditional approach using Theorem 1. We needed to numerically compute 1000 $\lambda(\boldsymbol{\eta}_{r})$ values and their corresponding derivatives in $\tau^{2}$ to approximate the LP distribution under the predictive approach, which took 15 seconds on a high computing server with 72 cores. Roughly one hour on this server was required to approximate all LP curves in Figure 2 via Algorithm 2. This process was computationally intensive because we approximated each posterior of $\boldsymbol{\eta}$ using Markov chain Monte Carlo with 1000 burnin iterations and 5000 retained draws, but it could have been expedited using analytical methods for posterior approximation. Nevertheless, Algorithm 2 is still much more efficient than naive simulation. On the same computing server, it took 9 and 19 hours to naively approximate the LP curve for this example under the conditional and predictive approaches, respectively. Our numerical results in Appendix C confirm the strong performance of Algorithm 2 in Bayesian settings and further demonstrate that analytical approximations based on Theorem 1 may perform poorly with moderate sample sizes when (i) informative priors are used or (ii) the posterior distributions used to construct the CI exhibit substantial skewness.

### v4

@@ Abstract
Estimation frameworks for statistical inference are preferred to hypothesis testing when quantifying uncertainty and precise estimation are more valuable than binary decisions about statistical significance. Study design for estimation-based investigations often uses precision criteria to select sample sizes that control the length of interval estimates with respect to a sampling distribution. In this paper, we formally define a distribution that characterizes the probability of obtaining a sufficiently narrow interval estimate as a function of the sample size. This distribution can be used to determine the smallest sample size needed to ensure an interval estimate is sufficiently narrow. We prove that this distribution is approximately normal in large-sample settings for many data generation processes. However, this approximate normality may not hold for studies with moderate sample sizes, particularly when incorporating prior information or obtaining asymmetric interval estimates. Thus, we also propose an efficient simulation-based approach to approximate the distribution for the sample size by estimating the sampling distribution of interval estimate lengths at only two sample sizes. Our methodology provides a unified framework for design with precision criteria in Bayesian and frequentist settings. We illustrate the broad applicability of this framework with several examples.
@@ 1 Introduction
Practitioners across disciplines design studies to collect data that meaningfully inform decision making in the presence of uncertainty. Different study objectives – such as assessing the plausibility of hypotheses or precisely estimating quantities of interest – motivate distinct approaches to study design. While many studies are designed with the goal of rejecting a null hypothesis, the null hypothesis significance testing (NHST) framework has been scrutinized for its association with the scientific replicability crisis (Open Science Collaboration, 2015). The American Statistical Association replied to these concerns by releasing an official statement on $p$-values (Wasserstein and Lazar, 2016) that challenged the statistical community to develop alternatives and extensions to the traditional hypothesis testing framework. This statement prompted various responses including a special issue of The American Statistician that discussed “moving to a world beyond p<0.05p<0.05(Wasserstein et al., 2019) along with calls to action from the President of the Institute of Mathematical Statistics (Meng, 2019) and from the Royal Statistical Society in Significance (Matthews, 2021).
@@
To accommodate Bayesian and frequentist approaches, we broadly introduce precision criteria for interval estimates. We henceforth use the term CI for statements that apply to both confidence intervals and credible intervals. Our methods accommodate various types of CIs – including frequentist confidence intervals as well as equal-tailed credible intervals or those defined using regions of highest posterior density. For now, we denote the length of the $100\times(1-\alpha)\%$ CI for the target of inference $\theta$ as $L_{1-\alpha}$. We later emphasize that this CI is a function of the data. Precision criteria generally select a sample size $n$ to ensure a large probability that $L_{1-\alpha}$ is sufficiently narrow with respect to the sampling distribution for the future data. The length probability criterion (LPC) was defined to formalize this objective in Bayesian settings (De Santis and Pacifico, 2004; Brutti et al., 2014). In a Bayesian or frequentist setting, the LPC could be applied to select the smallest sample size $n$ such that $\Pr(L_{1-\alpha}\leq l)\geq q$ for some target length $l$ and probability $q\in(0,1)$. This focus on controlling the length of CIs is not to the exclusion of the frequentist notion of coverage and its Bayesian analogues (Gustafson, 2012), which should approach the nominal level of $1-\alpha$ as $n\rightarrow\infty$ under standard regularity conditions (van der Vaart, 1998).
In this work, we define a distribution for the LPC with respect to the sample size $n$ and explore its theoretical properties. We call this distribution the length probability distribution (abbreivated as the LP distribution). The LP distribution is formally defined so that its $q$-quantile is the smallest sample size $n$ such that $L_{1-\alpha}$ will have length at most $l$ with probability at least $q$ (i.e., such that $\Pr(L_{1-\alpha}\leq l)\geq q$). This construct broadly accommodates Bayesian and frequentist CIs under both the conditional and predictive approaches to specify sampling distributions.
@@
In this paper, we propose two methods to estimate the LP distribution across statistical paradigms. The first method analytically approximates the LP distribution using asymptotic theory. The second method leverages this asymptotic theory to inform an efficient method to estimate the LP distribution based on simulation results conducted at only two sample sizes; this method is generally applicable and flexibly accommodates moderate sample size settings where asymptotic approximations may be unsuitable. While our design framework proposed here is theoretically intricate, its implementation is straightforward, promoting an economical and broadly useful approach to simulation-based sample size determination with precision criteria.
@@ 2 Preliminaries
Our design framework denotes data from a random, future sample as $\boldsymbol{W}^{{}_{(n)}}$, where the sample size $n$ is the number of independent sampling units (ISUs). The observed data are denoted by $\boldsymbol{w}^{{}_{(n)}}$. For non-clustered data, each observation is an ISU and in settings with clustered data, multiple dependent observations may comprise a single ISU. For each ISU, the sample $\boldsymbol{W}^{{}_{(n)}}$ may consist of response(s) $\boldsymbol{Y}$ along with a vector or matrix of additional covariates $\boldsymbol{X}$ that may encode information about treatment group assignment and/or assignment to ISUs. We assume that each ISU in $\boldsymbol{W}^{{}_{(n)}}$ is generated independently according to the model $f(\boldsymbol{w};\boldsymbol{\eta})$, where $\boldsymbol{\eta}$ denotes a vector of parameters. The target of inference $\theta$ is specified as a function $g(\cdot)$ of these parameters: $\theta=g(\boldsymbol{\eta})$. In an estimation framework, we aim to draw inference using a two-sided $100\times(1-\alpha)\%$ CI for $\theta$. We henceforth denote the length of the $100\times(1-\alpha)\%$ CI as $L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})$ for a future sample of $n$ ISUs; the length of the corresponding CI for an observed sample is $L_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}})$. This notation now emphasizes that the CI length and its sampling distribution are functions of the data and the confidence or credibility level $1-\alpha$.
@@
The collection of estimates $\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})\}_{r=1}^{m}$ in Line 5 of Algorithm 1 is used to estimate the sampling distribution of CI lengths under $\Psi$. The relevant $100\times(1-\alpha)\%$ CIs for $\theta$ can be estimated using any Bayesian or frequentist method that is based on likelihood functions. This flexibility accommodates the use of simple or complex interval estimates – such as approximate frequentist confidence intervals obtained using normal theory and sample summary statistics or Bayesian highest posterior density intervals based on computationally approximating the posterior distribution of $\theta$. While design methodology for likelihood-based inference is the focus of this paper, we discuss extensions for other interval estimates in Section 7. The probability of satisfying the length criterion can be estimated as
@@ 3 Approximate Normality of the Length Probability Distribution
We prove Theorem 1 in Appendix B.2 of the supplement, but we consider its practical implications here. Theorem 1 imposes constraints on $\Lambda(\boldsymbol{\eta}_{r})$ and $\tau^{2}$ to ensure the normal distribution that approximates the LP distribution for a given length $l$ has finite mean $\mu_{l}>0$ and standard deviation $\sigma_{l}>0$. These constraints impose implicit assumptions that the partial derivatives in (3) and (5) exist. For large sample sizes $n$ corresponding to small target lengths $l$, the values for $\mu_{l}$ and $\sigma_{l}$ from Theorem 1 can be used to determine quantiles of the LP distribution under the conditional approach as stated in Corollary 1.
@@ 4 A Proxy to the Sampling Distribution of CI Lengths
Flexible, simulation-based design methods that control the length of CIs require that we estimate the sampling distribution of CI lengths for various sample sizes $n$. We approximate this sampling distribution by generating data $\boldsymbol{w}^{{}_{(n)}}$ using the straightforward process in Algorithm 1. For theoretical development, we create a proxy for the sampling distribution of the logarithm of the CI lengths. These proxies substantiate the theory that underpins our proposed methodology, but they are not used in our design methods proposed in Section 5. In the hypothesis testing framework, proxies to sampling distributions of test statistics and posterior probabilities have also been used to expedite simulation-based design (Hagar and Stevens, 2024, 2025a, 2025c, 2025b).
@@ 6.2 Estimation of Illness Duration
We next assess the performance of our proposed methods in the context of estimating the duration of Mpox illness. The Mpox virus is predominantly spread via direct contact with the active lesions of an already infected individual (Kim et al., 2023). In previous clinical trials (Klein, 2024), Mpox illness duration was defined as the time in days between starting tecovirimat (a safe drug approved for use with smallpox) and experiencing the resolution of all Mpox lesions. For this example, we want to estimate the median of this distribution for Mpox illness duration. Our subsequent numerical results serve purely illustrative purposes, but we emphasize the value in precisely estimating quantities associated with illness durations to inform public health recommendations. This example also allows us to consider settings with independent observations, account for additional covariates, and accommodate censoring.
@@
We model the resolution time $T_{i}$ for participant $i$ using a Bayesian proportional hazards model. The maximum follow-up time is 4 weeks, so all participants have an administrative censoring time of $C_{i}=28$ days. The baseline “hazard” of experiencing lesion resolution is a piecewise constant function. This baseline hazard $h_{0}(t)$ changes with the week of observation such that $h_{0}(t)$ is $h_{1}$ if $0\leq t\leq 7$, $h_{2}$ if $7<t\leq 14$, $h_{3}$ if $14<t\leq 21$, and $h_{4}$ if $t>21$. We collectively refer to these constants as $\boldsymbol{h}=(h_{1},h_{2},h_{3},h_{4})$. Participant $i$’s hazard of experiencing lesion resolution in week $j$ is therefore $h_{j}\exp(\beta x_{i})$, where $\beta$ is a regression parameter. The to-be-observed time $Y_{i}$ for participant $i$ is the minimum of $T_{i}$ and $C_{i}$. At the design stage, we must also define a distribution for $\{X_{i}\}_{i=1}^{n}$. We suppose that $X\sim\mathcal{N}(0,\sigma_{x}^{2})$ for this example, and that the data from participants $i$ and $i^{\prime}$ are independent for $i\neq i^{\prime}$. To summarize, the parameters for our analysis model are $\boldsymbol{\eta}=(\boldsymbol{h},\beta)$, but the parameters $\boldsymbol{\eta}^{+}=(\boldsymbol{h},\beta,\sigma_{x})$ are used for data generation.
@@
Our target of inference $\theta$ used for study design is the median of the distribution for the resolution time $T$ when $x=0$. This median is an implicit function $g(\cdot)$ of the baseline hazards $\boldsymbol{h}$; it is the time at which the cumulative baseline hazard equals $\log(2)$. We aim to estimate a $100\times(1-\alpha)\%=95\%$ posterior highest density interval (HDI) for $\theta$ having length at most $l=2$ days with probability $q=0.8$. To demonstrate the broad applicability of our methods, we design this study under the conditional and predictive approaches. For the conditional approach, we use the parameter values $\boldsymbol{\eta}^{+}_{r}=((0.045,0.095,0.0325,0.01),0.5,0.25)$ informed by findings from Kim et al. (2023) for all simulation repetitions $r$. These parameter values are such that the true value of $\theta_{r}$ is 11 days and lesion resolution times are censored for roughly 28% of participants. Under the predictive approach, we assume that $\log(\boldsymbol{h}_{r})\sim\mathcal{N}(\log((0.045,0.095,0.0325,0.01)),0.07^{2}\times\mathbb{I}_{4})$ such that $\mathbb{I}_{4}$ is the $4\times 4$ identity matrix and $\beta_{r}\sim\mathcal{N}(0.5,0.07^{2})$; we note that small changes to these parameter values give rise to substantial changes in $\theta_{r}$ and the censoring proportions. We again assume that $\sigma_{x,r}=0.25$ for all repetitions under the predictive approach.
@@
To facilitate Bayesian analysis, we consider a diffuse and informative prior for $\boldsymbol{\eta}$ to explore how our methods perform. The diffuse prior is $\mathcal{N}(0,10^{2}\times\mathbb{I}_{4})$ for $\log(\boldsymbol{h})$ and $\mathcal{N}(0,10^{2})$ for $\beta$. The informative prior is $\mathcal{N}(\log((0.045,0.095,0.0325,0.01)),0.3^{2}\times\mathbb{I}_{4})$ for $\log(\boldsymbol{h})$ and $\mathcal{N}(0.5,0.3^{2})$ for $\beta$. This joint informative prior for $\log(\boldsymbol{h})$ and $\beta$ is relatively well specified for illustration. These choices result in four approach and prior combinations. For each combination, we analytically approximated the LP curve using the results from Theorem 1. For the predictive approach, this analytical approximation arises from a mixture of normal distributions that we obtained by applying Theorem 1 to 1000 simulated $\boldsymbol{\eta}^{+}_{r}$ values. We also approximated the LP curve using Algorithm 2 with $m=10^{4}$ simulation repetitions; to accommodate the predictive approach, the CI lengths were split into 10 bins based on the true $\Lambda(\boldsymbol{\eta}_{r})$ value at each sample size before constructing the linear approximations. We lastly approximated the LP curve by simulating the sampling distribution of CI lengths in increments of $n=10$. The resulting LP curve approximations are visualized in Figure 2.
@@
Figure 2: LP curves obtained using various estimation methods when the median of Mpox illness duration is the target of inference. The horizontal dotted line represents $q=0.8$.
@@
Algorithm 2 again yields more accurate LP curve approximation than the analytical approximation. The performance of the analytical approximation resulting from Theorem 1 is serviceable for the settings with the diffuse prior. Because the analytical approximation to the LP curve does not account for the prior distribution, it overestimates the required sample size when the relatively well-specified informative prior is used ($n=392$ vs. 357 for the conditional approach and $n=413$ vs. 376 for the predictive approach). Overall, the LP distributions under the predictive approach exhibit greater dispersion than those under the conditional approach. That is, we require less (more) data to attain the length criterion with smaller (larger) probabilities $q$. This greater dispersion in the LP distribution is a direct result of there being greater dispersion in the sampling distribution of CI lengths under the predictive approach.
@@
It took one second to approximate the LP distribution under the conditional approach using Theorem 1. The analytical approximation for this example is not instantaneous because we needed to numerically approximate the derivatives and Fisher information in (3) and (5). When approximating the Fisher information numerically, we can accommodate the straightforward censoring process in this example or more complex censoring mechanisms. We numerically computed 1000 $\lambda(\boldsymbol{\eta}_{r})$ values and their corresponding derivatives in $\tau^{2}$ to approximate the LP distribution under the predictive approach, which took 20 seconds on a high computing server with 72 cores. Roughly 14 minutes on this server was required to approximate all LP curves in Figure 2 via Algorithm 2. This process was computationally intensive because we approximated each posterior of $\boldsymbol{\eta}$ using Markov chain Monte Carlo with 1000 burnin iterations and 5000 retained draws, but it could have been expedited using analytical methods for posterior approximation. Nevertheless, Algorithm 2 is still much more efficient than naive simulation. On the same computing server, it took 3.5 and 5.5 hours to naively approximate the LP curve for this example under the conditional and predictive approaches, respectively. Our numerical results in Appendix C confirm the strong performance of Algorithm 2 in Bayesian settings and further demonstrate that analytical approximations based on Theorem 1 may perform poorly with moderate sample sizes when (i) informative priors are used or (ii) the posterior distributions used to construct the CI exhibit substantial skewness.